# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v1

@@ Abstract
It is well known that most foundations for Bishop’s constructive mathematics are incompatible with a classical predicative development of analysis as put forward by Weyl in his Das Kontinuum.
Here, we first recall how this incompatibility arises from the possibility, present in most constructive foundations, to define sets by quantifying over (the exponentiation of) functional relations. This possibility is not allowed in modern formulations of Weyl’s logical system.
Then, we argue how a possible way out is offered by foundations, such as the Minimalist Foundation, where exponentiation is limited to a primitive notion of function defined by $\lambda$-terms as in dependent type theory.
The price to pay is to renounce the so-called rule of unique choice identifying functional relations with $\lambda$-terms, and to number-theoretic choice principles, characteristic of foundations aimed to formalize Bishop’s constructive analysis. This restriction calls for a point-free constructive development of topology as advocated by P. Martin-Löf and G. Sambin with the introduction of Formal Topology.
Hence, we conclude that the Minimalist Foundation promises to be a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity provided that a point-free reformulation of classical analysis is viable.
@@ 1 Introduction
Another major reformulation of real analysis within a foundation weaker than the usual classical one was given by Bishop in [6]. While Bishop followed Brouwer in adopting intuitionistic logic, he refrained in his approach from including Brouwer’s continuity principles that are inconsistent with classical mathematics. Bishop’s aim was to develop constructive mathematics not in opposition to classical mathematics but as a generalization of the latter.
After the publication of Bishop’s book, there was a flourishing of foundations for constructive mathematics aimed at formally capturing Bishop’s approach. A relevant example in axiomatic set theory is Aczel’s constructive set theory CZF [1], and in type theory Martin-Löf Type Theory MLTT [38, 39] and the more recent Homotopy Type Theory HoTT  in [53] inspired by Voevodsky’s Univalent Foundations. In particular, HoTT  gained a lot of attention for its application to the formalization of synthetic homotopy theory and higher category theory. However, these systems are sometimes not even mutually compatible and this consideration led the second author in joint work with G. Sambin to conceive the Minimalist Foundation (for short MF) in [33], later fully worked out as a two-level system in [24]. The two-level structure of MF in [24], which includes an intensional level, an extensional level, and an interpretation of the latter into the first, was introduced to better accomplish the compatibility of MF with constructive and classical foundations including the mentioned ones (see [24, 9]) and internal theories of categorical universes such as that of elementary topoi in [22, 23].
@@
Here we first discuss the source of this incompatibility. The reason lies in their possibility of defining sets by quantifying over (the exponentiation of) functional relations. Such a possibility is not allowed in modern formulations of Weyl’s system [5, 4, 3, 16].
Then, we argue that it is not necessary to sacrifice exponentiation altogether to reconcile Bishop’s mathematics with Weyl’s approach if we adopt a foundation like the Minimalist Foundation and this for a number of reasons.
@@
Second, the classical version of MF is equi-consistent, as shown in [27, 44], with the original constructive MF, whose proof-theoretic strength is strictly predicative a’ la Feferman being interpretable in first-order Martin-Löf’s type theory [24] and directly in Feferman’s $\widehat{ID_{1}}$ (see [20]).
@@
The price to be paid in MF  is to renounce the so-called rule of unique choice identifying functional relations with $\lambda$-terms. In this way exponentiation of type-theoretic functions does not entail exponentiation of functional relations as it happens instead in Martin-Löf’s type theory and in Homotopy Type Theory, whose logics are also sorted on a dependent type theory. Accordingly, dropping the rule of unique choice entails dropping all the number-theoretic choice principles characteristic of Bishop’s mathematics as well as the fact that real numbers a’ la Dedekind or Cauchy form a set.
As a consequence, these limitations call for a point-free constructive development of topology as advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45] to avoid the use of Brouwer’s Fan theorem as explained in [34]. Evidence that such an approach can be successfully developed is provided by the point-free results in [7, 41, 42, 7, 8, 21, 47].
Given that the point-free topology of Dedekind real numbers and that of Cantor space are definable in MF, if a point-free reformulation of classical analysis is fully viable, then MF  would become a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity.
In the future, besides determining the exact proof-theoretic strength of MF, which is currently an open problem, we intend to investigate whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] is still predicative, or even still equi-consistent with their classical counterparts as established for MF in [27, 44]. A positive answer to one of these questions would provide a foundation more expressive than MF  and capable of formalizing a whole development of point-free topology, including for example the results in [36, 15, 10, 47], while still maintaining compatibility with both Bishop’s constructive mathematics and Weyl’s classical predicative mathematics.
@@ 2 Weyl’s Classical Predicativity
Das Kontinuum [55, 56] can be regarded as the first systematic attempt to develop predicative mathematics. The main focus of the book is the foundation of analysis and, moreover, the reconstruction of many fundamental results in this area without resorting to impredicative principles. Examples of theorems shown by Weyl in his framework include the Fundamental Theorem of the Calculus, Uniform continuity for continuous functions over a closed interval and existence of a maximum and a minimum for them as well as sequential versions of the Heine-Borel Theorem and the Least Upper Bound principle. Therefore, the main merit of Das Kontinuum was to show that predicative mathematics was a viable project. Feferman has shown that the above mentioned results can be formalized in the subsystem of second-order arithmetic known as ACA0 and further evidence in support of this claim has been provided by the Reverse Mathematics program [48]. In this section, we review some basic tenets of Weyl’s approach in relation to the Minimalist Foundation, which will be introduced later in Section 6.
Weyl acknowledged - in agreement with Poincaré- the absolutely primitive character of the natural numbers and the associated notion of iteration. According to Weyl, every attempt to understand natural numbers in terms of more fundamental notions – as it happens in Zermelo’s set theory or in the logicist program of developing arithmetic on the basis of purely logical notions– was doomed to failure because it necessarily presupposes this very notion:
@@
Weyl’s conception of set bears some resemblance to the way sets are understood in Martin-Löf’s type theory (MLTT) and also in both levels of the Minimalist Foundation (MF), which will be presented in the next sections. In such theories, a set is determined by rules for forming its elements equipped with an induction principle to prove properties on it and a recursive definition to define a function by recursion from it to another set, see [39].
@@
In particular, it follows from this quotation that the power of a set is not a set according to Weyl. Thus we cannot define a new set by quantifying over the powerset to which it should belong. The subtle point is that this does not mean that the power of a set does not exist, but simply that it is not a set but a collection that is not inductively generated. In the Minimalist Foundation, this distinction between sets and collections is explicitly assumed. Accordingly, in Weyl’s work collections exceed the power-collection of a set, as shown in [4, 3, 5], but they are not allowed to enter into the definition of new sets.
@@
Both in Das Kontinuum and in the Minimalist Foundation real numbers provide an example of a collection which is not a set. In particular, Weyl introduced real numbers as Dedekind cuts of rationals and are formalized as quadruples of natural numbers. The fact that the resulting collection $\mathbb{R}$ of real numbers is not extensionally determinate in Weyl’s terminology, means that we cannot define new sets by quantification over $\mathbb{R}$, in sharp contrast with the case of natural numbers. This implies, for instance, that the least upper bound principle for sets of real numbers is not predicatively provable, while the principle is provable for sequences of real numbers, since the latter do not involve quantification over $\mathbb{R}$.
Finally, we want to point out that Weyl in Das Kontinuum places a significant emphasis on his notion of function and contrasts it with the set-theoretical definition. To this purpose, Avron writes in [5]:
‘Weyl totally rejected the modern view of a function as any set of pairs that satisfies the functionality condition. He insists that (like sets) functions can be determined only by a rule’.
@@
With reference to this, it might be intriguing to attribute to Weyl a kind of computational understanding of what a function is. Indeed, Weyl works in a classical framework and assumes that functions are extensional objects just as sets in his system. However, Weyl’s quotation is in line with the fact that in the Minimalist Foundation we distinguish between functions as functional relations and functions as primitive operations defined by $\lambda$-terms. The distinct character of these two notions of functions is a peculiarity of MF among foundations for constructive predicative mathematics. While it might be argued that Weyl foreshadowed this distinction, nevertheless his technical treatment of the notion of function is not exactly the same. Therefore, we can only highlight the significance of the insight of viewing functions as rule-based objects. Moreover, this resembles the notion of Brouwer lawlike sequence, which we will identify with a $\lambda$-term in the Minimalist foundation (see [46]).
@@
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion, since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been offered by Feferman in several works [16, 17, 18]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [48]. However, Feferman’s proposal has been criticized by Adams and Luo in [4, 3] and more recently by Avron in [5]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than what thought by Feferman, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
@@
The systems introduced in these papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. But they differ between them and with MF for what regards exponentiation of functions. Indeed, Adams and Luo’s system in [4, 3] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Then, Avron in [5] restricts the possibility to form exponentials by assuming that only certain types might serve as codomains. However, in his system we cannot define a set by quantifying over functions. Instead, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [1] and Martin-Löf Type Theory MLTT  [39]. More recently, an extension of the latter called Homotopy Type Theory HoTT [53] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity a lá Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or of the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.1 Aczel’s Constructive Zermelo-Fraenkel set theory
cannot be added to CZF while preserving its constructive character:
@@
CZF + AC derives the law of excluded middle for $\Delta_{0}$-formulas.
@@ 3.2 Martin-Löf’s type theory
Types of MLTT are introduced by means of four rules: formation, introduction, elimination and computation. In particular, the constructors of a given type are specified by the introduction rules, while the elimination rule states simultaneously a principle of induction at the level of types and of definition by recursion at the level of terms with the aid of the computation rules. The predicative nature of the theory is related to the fact that all the types are inductively generated by the relevant rules. Two fundamental aspects of MLTT are the presence of a primitive notion of function, called here $\lambda$-function to disambiguate it from the notion of functions as functional relation, and the identification between propositions and types. $\lambda$-functions are represented as $\lambda$-terms and this differentiates this foundational theory from other theories such as CZF. The identification between propositions and types makes the constructive logic of MLTT stronger than usual intuitionistic logic. This additional strength is visible in the case of the existential quantifier that is identified with the dependent sums type $\Sigma$. The elimination rule for this type is enough strong to allow for the definition of projection operators  [^2] We recall that the canonical elements of $\Sigma$-types given by the introduction rule are pairs.. These projections play a fundamental role in establishing the following fact:
@@
is derivable in MLTT.
@@
Due to the absence of quotient sets in MLTT, real numbers, both à la Dedekind or à la Cauchy, are formally represented there using setoids whose elements are then indexed by a set. A setoid [40], called set by Bishop in ([6], p. 13), is a pair $(A,=_{A})$, consisting of a set called support equipped with a suitable equivalence relation $a=_{A}a^{\prime}\ [a,a^{\prime}:A]$ used to equate the elements in the same equivalence class.
For example, Bishop reals in [6] can be represented as the setoid whose support is given by the set of Cauchy regular sequences from the set of natural numbers to that of rational numbers equipped with the equivalence relation $=_{\mathcal{R_{C}}}$ as in [6], pp.15-16. Bishop defined them by using the notion of operation which we can identify in MLTT  with the notion of $\lambda$-function, i.e. with a $\lambda$-term of MLTT. Moreover, for what we have said about the notion of function in MLTT, Cauchy reals defined as Cauchy functional relations can be identified in MLTT with Bishop reals. For unifying the various concepts, we name them Cauchy sequences a’ la Bishop.
@@
A Dedekind cut on the rationals is a pair of propositional functions $L,U:\mathbb{Q}\rightarrow\mathcal{U}_{k}$ for some universe $U_{k}$ that satisfy the usual conditions of being inhabited, disjoint, open, monotone, and located, all defined as in [35].
@@
Again thanks to the validity in MLTT of choice principles on natural numbers and thanks to the exponentiation of $\lambda$-terms by means of the dependent product type $\Pi_{(x:A)}B(x)$ of any family of types $B(x)(x:A)$, in MLTT each real number à la Dedekind in the form of a Dedekind cut corresponds bijectively to a Bishop real and hence to a Dedekin cut in $\mathbb{R}_{D_{0}}$.
@@
In MLTT, Dedekind reals $\mathbb{R}_{D_{0}}$ are in bijective correspondence with $\mathbb{R}_{D_{k}}$ for any natural number $k$.
@@
The previous proposition shows that some instances of resizing can be derived from choice principles constructively. In particular, this can be regarded as the fact that special cases of Russell’s Reducibility Axiom can be derived in the presence of choice principles provided that the logic is constructive.

### v2

@@ Abstract
It is well known that most constructive and predicative foundations aiming to develop Bishop’s constructive analysis are incompatible with a classical predicative development of analysis as Weyl put forward in his Das Kontinuum.
Here, we show how this incompatibility arises from the possibility to define sets by quantifying over (the exponentiation of) functional relations. Such a possibility is present in most constructive foundations but it is not allowed in modern reformulations of Weyl’s logical system. In particular, we argue how in Aczel’s Constructive Set Theory, Martin-Löf’s type theory and Homotopy Type Theory, the incompatibility with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
Then, we argue how a possible way out is offered by constructive foundations, such as the Minimalist Foundation, where exponentiation is limited to functions defined as $\lambda$-terms of (dependent) type theory.
The price to pay is to renounce number-theoretic choice principles, including the rule of unique choice, typical of most foundations formalizing Bishop’s constructive mathematics. This restriction calls for a point-free constructive development of topology as advocated by P. Martin-Löf and G. Sambin with the introduction of Formal Topology.
We then conclude that the Minimalist Foundation promises to be a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity provided that a point-free constructive reformulation of analysis is viable.
@@ 1 Introduction
Another major reformulation of real analysis within a foundation weaker than the usual classical one was given by Bishop in [6]. While Bishop followed Brouwer in adopting intuitionistic logic, he refrained in his approach from including Brouwer’s continuity principles that are inconsistent with classical mathematics. Bishop aimed to develop constructive mathematics not in opposition to classical mathematics but as a generalization of the latter.
After the publication of Bishop’s book, there was a flourishing of foundations for constructive mathematics aimed at formally capturing Bishop’s approach. A relevant example in axiomatic set theory is Aczel’s constructive set theory CZF [1], and in type theory Martin-Löf Type Theory MLTT [39, 40] and the more recent Homotopy Type Theory HoTT  in [53] inspired by Voevodsky’s Univalent Foundations. In particular, HoTT  gained a lot of attention for its application to the formalization of synthetic homotopy theory and higher category theory. However, these systems are sometimes not even mutually compatible and this consideration led the second author in joint work with G. Sambin to conceive the Minimalist Foundation (for short MF) in [34], later fully developed as a two-level system in [24]. The two-level structure of MF in [24], which includes an intensional level, an extensional level, and an interpretation of the latter into the first, was introduced to better achieve the compatibility of MF with constructive and classical foundations including those mentioned (see [24, 9]) as well as internal theories of categorical universes such as that of elementary topoi in [22, 23].
@@
Here we first discuss the source of this incompatibility. The reason lies in their possibility of defining sets by quantifying over (the exponentiation of) functional relations. Such a possibility is not allowed in modern formulations of Weyl’s system [5, 4, 3, 16]. More formally, we show that the incompatibility of CZF, MLTT and HoTT with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
Then, we argue that it is not necessary to sacrifice exponentiation altogether to reconcile Bishop’s mathematics with Weyl’s approach if we adopt a foundation like the Minimalist Foundation and this for various reasons.
@@
Second, the classical version of MF is equiconsistent, as shown in [27, 33], with the original constructive MF, whose proof-theoretic strength is strictly predicative à la Feferman being interpretable in first-order Martin-Löf’s type theory with one universe [24] and directly in Feferman’s $\widehat{ID_{1}}$ (see [20]).
@@
The price to be paid in MF  is to renounce the so-called rule of unique choice identifying functional relations with $\lambda$-terms. In this way, exponentiation of type-theoretic functions does not entail exponentiation of functional relations as it happens instead in Martin-Löf’s type theory and in Homotopy Type Theory, whose logics are also sorted on a dependent type theory. Accordingly, dropping the rule of unique choice entails dropping all the number-theoretic choice principles characteristic of Bishop’s mathematics as well as the fact that real numbers à la Dedekind or Cauchy form a set.
As a consequence, these limitations call for a point-free constructive development of topology as advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45] to avoid the use of Brouwer’s Fan theorem as explained in [35]. Evidence that such an approach can be successfully developed is provided by the point-free results in [7, 42, 43, 8, 21, 47].
Given that the point-free topology of Dedekind real numbers and that of Cantor space are definable in MF, if a point-free reformulation of classical analysis were fully viable, then MF  would become a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity.
In the future, besides determining the exact proof-theoretic strength of MF, which is currently an open problem, we intend to investigate whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterparts as established for MF in [27, 33], or at least these classical counterparts are still predicative. A positive answer to one of these questions would provide more expressive foundations than MF capable of formalizing a whole development of point-free topology, including for example the results in [37, 15, 10, 47], while still keeping the compatibility with both Bishop’s constructive mathematics and Weyl’s classical predicative mathematics.
@@ 2 Weyl’s Classical Predicativity
Das Kontinuum [55, 56] can be regarded as the first systematic attempt to develop predicative mathematics. The book’s main focus is the foundation of analysis and the reconstruction of many fundamental results in this area without resorting to impredicative principles. Examples of theorems shown by Weyl in his framework include sequential versions of the Heine-Borel Theorem and the Least Upper Bound principle, Uniform continuity for continuous functions over a closed interval, the existence of a maximum and a minimum for them as well as the Fundamental Theorem of the Calculus.
Therefore, the main merit of Das Kontinuum was to show that predicative mathematics was a viable project. Feferman has shown that the above mentioned results can be formalized in the subsystem of second-order arithmetic known as ACA0 and further evidence in support of this claim has been provided by the Reverse Mathematics program [48]. In this section, we review some basic tenets of Weyl’s approach by relating them to the Minimalist Foundation, which will be introduced later in Section 6.
Weyl acknowledged - in agreement with Poincaré- the primitive character of the natural numbers and the associated notion of iteration. According to Weyl, every attempt to understand natural numbers in terms of more fundamental notions – as it happens in Zermelo’s set theory or in the logicist program of developing arithmetic from purely logical notions– was doomed to failure because it necessarily presupposes this very notion:
@@
Weyl’s conception of set bears some resemblance to the way sets are understood in Martin-Löf’s type theory (MLTT) and also in both levels of the Minimalist Foundation (MF), which will be presented in the next sections. In such theories, a set is determined by rules for forming its elements, equipped with an elimination rule that allows proving properties by induction over it and defining functions by recursion from it to another set, see [40].
@@
In particular, it follows from this quotation that the power of a set is not a set according to Weyl. Thus we cannot define a new set by quantifying over the powerset to which it should belong. The subtle point is that this does not mean that the power of a set does not exist, but simply that it is not a set but a collection that is not inductively generated.
In the Minimalist Foundation, this distinction between sets and collections is explicitly assumed. Accordingly, in Weyl’s work collections exceed the power-collection of a set, as shown in [4, 3, 5], but they are not allowed to enter into the definition of new sets.
@@
Both in Das Kontinuum and in the Minimalist Foundation real numbers provide an example of a collection which is not a set. In particular, Weyl introduced real numbers as Dedekind cuts of rationals that are formalized as quadruples of natural numbers. The fact that the resulting collection $\mathbb{R}$ of real numbers is not extensionally determinate in Weyl’s terminology, means that we cannot define new sets by quantification over $\mathbb{R}$, in sharp contrast with the case of natural numbers. This implies, for instance, that the least upper bound principle for sets of real numbers is not predicatively provable, while the principle is provable for sequences of real numbers, since the latter do not involve quantification over $\mathbb{R}$.
Finally, we want to point out that Weyl in Das Kontinuum places a significant emphasis on his notion of function and contrasts it with the traditional set-theoretical definition (see [56], pp.33-34). According to Avron’s analysis in [5], for Weyl, a function is always determined by a rule. In other words, a function cannot be given by an arbitrary relation satisfying the functionality condition, as in set theory; rather, there are syntactic constraints on the definability of the relation that must be respected.
In this regard, it might be intriguing to relate Weyl’s conception of function as a rule-based object to a computational understanding of functions.
@@
Weyl’s insights about functions as rule-based objects align, in some way, with the fact that the Minimalist Foundation distinguishes the notion of function as functional relation from that of function as a primitive operation defined by a $\lambda$-term, resembling the Brouwer’s notion of lawlike sequence (see [46]). The distinct character of these two notions of function is a peculiarity of MF among foundations for constructive predicative mathematics. While it might be argued that Weyl foreshadowed this distinction, nevertheless his technical treatment of the notion of function is not exactly the same.
@@
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been carried on by Feferman in several works [16, 17, 18]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [48]. But such a reconstruction has been criticized by Adams and Luo in [4, 3], and more recently by Avron in [5]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than Feferman’s reformulation of Weyl’s system, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
@@
The systems introduced in these papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. They all differ in the way exponentiation of functions is treated. Indeed, Adams and Luo’s system in [4, 3] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Instead, Avron in [5] restricts the possibility of forming exponentials by assuming that only certain types might serve as codomains by not allowing to define sets by quantification over functions. On the contrary, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [1] and Martin-Löf Type Theory MLTT  [40]. More recently, an extension of the latter called Homotopy Type Theory HoTT [53] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity a lá Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.1 Aczel’s Constructive Zermelo-Fraenkel set theory
where $A$ and $B$ are sets and $R$ any relation on them, cannot be added to CZF while preserving its constructive character:
@@
CZF + AC derives the law of excluded middle $\phi\vee\neg\phi$ for $\Delta_{0}$-formulas $\phi$.
@@ 3.2 Martin-Löf’s type theory
Types of MLTT are introduced through the following kinds of rules: formation, introduction, elimination, and computation. In particular, the constructors of a given type are specified by the introduction rules, while the elimination rule states simultaneously a principle of induction at the level of types and of definition by recursion at the level of terms with the aid of the computation rules. The predicative nature of the theory is related to the fact that all the types, also called sets, are inductively generated by the relevant rules. Two fundamental aspects of MLTT are the presence of a primitive notion of function, called here $\lambda$-function to disambiguate it from the notion of function as functional relation, and the identification between propositions and types. $\lambda$-functions are represented as $\lambda$-terms and this differentiates this foundational theory from other theories such as CZF. The identification between propositions and types makes the constructive logic of MLTT stronger than usual intuitionistic logic. This additional strength is visible in the case of the existential quantifier that is identified with the dependent sum type $\Sigma$. The elimination rule for this type is enough strong to allow for the definition of projection operators [^2] We recall that the canonical elements of $\Sigma$-types given by the introduction rule are pairs.. These projections play a fundamental role in establishing the following fact:
@@
where $A$ and $B$ are sets (i.e. types) and $R$ any relation on them, is derivable in MLTT.
@@
Due to the absence of quotient sets in MLTT, real numbers, both à la Dedekind or à la Cauchy, are formally represented there using setoids whose elements are then indexed by a set. A setoid [41], called set by Bishop in ([6], p. 13), is a pair $(A,=_{A})$, consisting of a set called support equipped with a suitable equivalence relation $a=_{A}a^{\prime}\ [a,a^{\prime}:A]$
For example, Bishop reals in [6] can be represented as the setoid whose support is given by the set of Cauchy regular sequences of rational numbers equipped with the equivalence relation $=_{\mathcal{R_{C}}}$ as in ([6], pp.15-16). Bishop defined them by using the notion of operation which we can identify in MLTT  with the notion of $\lambda$-function, i.e. with a $\lambda$-term of MLTT. Moreover, for what we have said about the notion of function in MLTT, Cauchy reals defined as Cauchy functional relations can be identified in MLTT with Bishop reals. For unifying the various concepts, we name them Cauchy sequences à la Bishop.
@@
A Dedekind cut on the rationals is a pair $(L,U)$ of subsets $L,U\subseteq\mathbb{Q}$ that satisfy the usual conditions of being inhabited, disjoint, open, monotone, and located, all defined as in [36].
Notice that the subsets $L,U$ are defined as elements of the setoid of propositional functions $L,U:\mathbb{Q}\rightarrow\mathcal{U}_{k}$ for some universe $\mathcal{U}_{k}$ quotiented under equiprovability. Therefore, ‘$a\ \varepsilon\ L$’ means that there exists a proof-term of $L(a)$.
@@
Again thanks to the validity in MLTT of choice principles on natural numbers and thanks to the exponentiation of $\lambda$-terms by means of the dependent product type $\Pi_{(x:A)}B(x)$ of any family of types $B(x)(x:A)$, in MLTT each real number à la Dedekind in the form of a Dedekind cut corresponds bijectively to a Bishop real and hence to a Dedekind cut in $\mathbb{R}_{D_{0}}$.
@@
In MLTT, the set $\mathbb{R}_{D_{0}}$ of Dedekind reals is in bijective correspondence with the set of Dedekind reals $\mathbb{R}_{D_{k}}$ for any natural number $k$.
@@
The previous proposition resembles a special case of Russell’s Reducibility Axiom which constructively requires a choice principle to be derived. If the law of excluded middle were valid the full Reducibility axiom would follow as shown in section 5.